# On Mean Field Games in Infinite Dimension

### v2

@@ 7.1. An economic model
with $\theta>0$ and $\pi:(0,\infty)^{2}\rightarrow\mathbb{R}$, $(k,\mathcal{K})\mapsto\pi(k,\mathcal{K})$ is nondecreasing in $k$ and nonincreasing in $\mathcal{K}$ [^6] In [9], the case $\theta=1$ and $\pi(k,\mathcal{K})=\mathcal{K}^{-\beta}k$, where $\beta>0$, is considered.
@@ Appendix A Compactness in the space of probability measures
By [39, Ch. VI, Th. 2.2] [^8] In [39, Ch. VI, Th. 2.2] it is stated that the second condition alone guarantees relative compactness of $S$. This is clearly not true, as can be seen by simply taking as $S$ a non compact sequence concentrated on the line generated by $\mathbf{e}_{1}$. However the proof of this theorem there remains valid if we add the assumption of uniform boundedness of second moments., the set $S$ is tight. The claim follows using [2, Prop. 7.1.5].
@@ Appendix B Classical and strong solutions to Kolmogorov equations
In this section, we define classical, mild, and $\mathcal{K}$-strong solutions of backward Kolmogorov equations and recall an approximation result used in Section 6. The material is mainly taken from [21, Section B.7] and is presented, for the reader’s convenience, in a simplified form needed here. We first consider the following terminal value problem (Kolmogorov equation) in the Hilbert space $H$:
@@
Arguing as in Theorem 2.7 of [18], it can be proved that $D(L_{0})$, endowed with the above norm, is a Banach space [^9] The definition of the domain of the operator studied in this paper is slightly different, but the arguments of the mentioned reference can be adapted easily. We also notice that the Banach space structure is not essential for our purposes, even if it simplifies the notation.. We now recall a notion of classical solution to (B.1) (see [21, Definition B.82]) [^10] We also observe that there are other definitions of classical solutions in the literature, see e.g. Section 6.2 of [19, Section 6.2], [27, Definition 4.6], [28, Definition 4.1]. We refer interested readers to [21, Section B.7.1] for a detailed overview on that..
@@
To introduce a notion of a strong solution, called $\mathcal{K}-$strong solution, we first recall a special case of the definition of $\mathcal{K}-$convergence which is needed in this paper (see [21, Definition B.84]).
@@
The definition of a $\mathcal{K}$-strong solution to (B.1) provided below is a special case of a more general definition that can be found in [21, Definition B.85].
@@
We end this section with an approximation result which is a stronger version of [21, Theorem B.95(i)].
@@
We can apply [21, Theorem B.95 (i)] (in the case when, in the notation there, $m=0$, $\eta(t)=t^{-\gamma}$), to conclude that $u$ is also a $\mathcal{K}$-strong solution of the same equation. Moreover, from the proof of [21, Theorem B.95 (i)] one can easily see that the approximating data $\varphi_{n}$, $f_{n}$, and consequently the approximating solutions $u_{n}$, can be chosen so that $\varphi_{n}\in D(L_{0})$ and $f_{n},\partial_{t}\bar{v}_{n}\in C_{b}([0,T]\times H)$ for all $n\in\mathbb{N}$. As a consequence, we get that $u_{n}\in\mathcal{D}_{T}$ for all $n\in\mathbb{N}$.

### v3

@@ 7.1. An economic model
with $\theta>0$ and $\pi:(0,\infty)^{2}\rightarrow\mathbb{R}$, $(k,\mathcal{K})\mapsto\pi(k,\mathcal{K})$ is nondecreasing in $k$ and nonincreasing in $\mathcal{K}$. [^6] In [10], the case $\theta=1$ and $\pi(k,\mathcal{K})=\mathcal{K}^{-\beta}k$, where $\beta>0$, is considered.
@@ Appendix A Compactness in the space of probability measures
By [40, Ch. VI, Th. 2.2] [^8] In [40, Ch. VI, Th. 2.2] it is stated that the second condition alone guarantees relative compactness of $S$. This is clearly not true, as can be seen by simply taking as $S$ a non compact sequence concentrated on the line generated by $\mathbf{e}_{1}$. However the proof of this theorem there remains valid if we add the assumption of uniform boundedness of second moments., the set $S$ is tight. The claim follows using [2, Prop. 7.1.5].
@@ Appendix B Classical and strong solutions to Kolmogorov equations
In this section, we define classical, mild, and $\mathcal{K}$-strong solutions of backward Kolmogorov equations and recall an approximation result used in Section 6. The material is mainly taken from [22, Section B.7] and is presented, for the reader’s convenience, in a simplified form needed here. We first consider the following terminal value problem (Kolmogorov equation) in the Hilbert space $H$:
@@
Arguing as in Theorem 2.7 of [19], it can be proved that $D(L_{0})$, endowed with the above norm, is a Banach space [^9] The definition of the domain of the operator studied in this paper is slightly different, but the arguments of the mentioned reference can be adapted easily. We also notice that the Banach space structure is not essential for our purposes, even if it simplifies the notation.. We now recall a notion of classical solution to (B.1) (see [22, Definition B.82]) [^10] We also observe that there are other definitions of classical solutions in the literature, see e.g. Section 6.2 of [20, Section 6.2], [28, Definition 4.6], [29, Definition 4.1]. We refer interested readers to [22, Section B.7.1] for a detailed overview on that..
@@
To introduce a notion of a strong solution, called $\mathcal{K}$-strong solution, we first recall a special case of the definition of $\mathcal{K}$-convergence which is needed in this paper (see [22, Definition B.84]).
@@
The definition of a $\mathcal{K}$-strong solution to (B.1) provided below is a special case of a more general definition that can be found in [22, Definition B.85].
@@
We end this section with an approximation result which is a stronger version of [22, Theorem B.95(i)].
@@
We can apply [22, Theorem B.95 (i)] (in the case when, in the notation there, $m=0$, $\eta(t)=t^{-\gamma}$), to conclude that $u$ is also a $\mathcal{K}$-strong solution of the same equation. Moreover, from the proof of [22, Theorem B.95 (i)] one can easily see that the approximating data $\varphi_{n}$, $f_{n}$, and consequently the approximating solutions $u_{n}$, can be chosen so that $\varphi_{n}\in D(L_{0})$ and $f_{n},\partial_{t}\bar{v}_{n}\in C_{b}([0,T]\times H)$ for all $n\in\mathbb{N}$. As a consequence, we get that $u_{n}\in\mathcal{D}_{T}$ for all $n\in\mathbb{N}$.